# Causal Learning for Trustworthy Recommender Systems: A Survey

### v1

@@ 2.2 Challenges in Representation Learning
Representation learning plays a fundamental role in bridging the stages of data preparation and recommendation generation. Although numerous advanced representation learning methods have been developed in RS, such as Recurrent Neural Networks (RNN), Graph Neural Networks (GNN) Ma et al. (2020) and transformer-based models Sun et al. (2019), they commonly overlook potential risks regarding fairness and robustness.
@@ 2.2.1 Fairness Challenges in Representation Learning
A major fairness challenge during representation learning is how to address conformity bias Zheng et al. (2021). This bias is closely related to popularity bias, but it specifically focuses on the perspective of user intentions. Generally speaking, a user-item interaction is usually caused by two factors: users’ genuine interests in the item and their conformity to the item’s popularity. An unbiased model should focus on learning users’ true interests rather than being misled by their conformity. However, it is intractable to clearly separate conformity intention from user representation, especially relying solely on data-driven approaches. To tackle this challenge, Zheng et al. Zheng et al. (2021) introduced a causality-driven framework, dubbed DICE, to learn representations with disentangled intentions. It presents a simple SCM model to depict the causal relationships between user implicit intentions and the resulting interactions. Leveraging the principle of the colliding effect Pearl (2009), DICE is adept at learning disentangled representations from cause-specific data.
@@ 2.2.2 Robustness Challenges in Representation Learning
Apart from explicit robustness risks in data, underlying noises can also significantly hinder the effectiveness of representation learning. Here, we mainly focus on two types of noises, i.e., spurious correlations He et al. (2023) and noisy feedback Zhang et al. (2021b).
Spurious correlation refers to a correlation that is statistically strong but in fact is causally incorrect. To clarify this concept, consider a straightforward example. It is a well-established fact that an item’s quality influences both its price and the users’ preference for it. Commonly, one can observe that higher-quality items are more expensive and more appealing to users. However, if the factor of quality is neglected, an RS may falsely imply that user preference increases along with the price. Such spurious correlations can be largely misleading. Unfortunately, data-driven methods, which typically rely on correlations or associations among samples, can hardly identify and eliminate such noises. In contrast, causality-oriented methods excel in tackling this issue by modeling and adjusting underlying causal relationships. For instance, recognizing that spurious correlations often stem from confounders, a study by Zhu et al. (2022) performs structure adjustment to mitigate corresponding confounder effects. Another representative research by Mu et al. (2022) introduces a counterfactual-based method to bolster model robustness against spurious correlations by generating high-quality counterfactual interactions.
Noisy feedback manifests in two ways: either as positive interactions made by users who in the end turned out to be unsatisfied Wang et al. (2021a, c), or as delayed interactions that are actually satisfied Zhang et al. (2021b). Such noisy feedback eventually limits the ability of an RS to capture users’ genuine preferences. To defend against this type of noise, Zhang et al. Zhang et al. (2021b) devised a counterfactual modifier in a reinforcement learning framework for streaming recommendation. It adopts importance sampling Yasui et al. (2020) and modifies the weight of reward in a counterfactual world to alleviate the impact of delayed feedback.
@@ 2.3 Challenges in Recommendation Generation
Upon obtaining informative representations from input data, RS models are devised to generate accurate suggestions that are tailored to individual preferences. Despite the promising performance that modern RS models Lin et al. (2022) achieved, there still exist several challenges regarding fairness and explainability due to their intrinsic mechanisms.
@@ 2.3.1 Fairness Challenges in Recommendation Generation
Additionally, feedback loops in RS often reinforce users’ existing preferences, creating an echo chamber that isolates them from diverse contents, known as the filter bubble bias Gao et al. (2024). To counter this challenge, a framework based on counterfactual inference Wang et al. (2022b) has been developed. This approach allows users to determine whether actions are needed to modify user embeddings through counterfactual modification. By doing so, it facilitates access to a wider and more diverse range of contents, effectively helping users to break out of filter bubbles.
2.3.2 Explainability Challenges in Recommendation Generation
The advancement of machine learning-based, especially deep learning-based techniques, has greatly enhanced the capability of RS models over the past decades. However, the inherent drawback of lacking interpretability has garnered significant concerns in recent years. With the introduction of complex black-box modules, it is challenging to explain how and why the recommendations are generated. This limitation not only impedes researchers in comprehensively understanding and improving models, but also undermines stakeholders’ trust in RS, especially in some safety-critical applications. Therefore, effective explainers are required for providing insightful explanations for both recommendation mechanisms and results.
Despite several attempts Zhang and Chen (2020) to generate explanations based on tags Vig et al. (2009) or visual attentions Chen et al. (2019), they are not convincing enough and often struggle to explain the mechanism of RS models. To bridge this gap, causality has been explored in RS to generate insightful explanations. For instance, a reinforcement learning model Xian et al. (2019) is designed to discover reasonable causal paths as explanations for how recommendations are generated. However, such path-based explanations also raise concerns regarding privacy leakage. Another representative work Tan et al. (2021) introduces a model-agnostic explainer based on counterfactual inference. This explainer aims to find straightforward explanations by searching for a set of changes on feature vectors, which is the least complex but strong enough to alter recommendation outcomes.
@@ 2.4 Challenges in Evaluation
Moving beyond accuracy, TRS encompass a variety of trustworthiness objectives, posing the necessity of specialized evaluation metrics. In particular, fairness metrics Li et al. (2023c) typically measure group or individual disparities. For instance, Wang et al. Wang et al. (2022b) employed an isolation index to measure the segregation across different user groups and to assess the severity of filter bubbles. In evaluating robustness, most works Li et al. (2023a) test RS accuracy under varying levels of robustness challenges. In addition, Li et al. Li et al. (2023b) proposed a novel metric to assess the balance of learned propensity weights when dealing with selection bias. In evaluating explainability, user reviews Yu et al. (2023) are usually collected to assess the consistency with generated explanations. Moreover, Tan et al. Tan et al. (2021) introduced two novel metrics, based on the concepts of necessity and sufficiency, to provide a thorough evaluation of explanation quality.
However, it should be pointed out that only a few attempts Sato et al. (2019) have been made to devise evaluations from a causal learning perspective. It is an important yet underexplored area in CTRS and it deserves wider research attention in the future. Moreover, we keep in mind that the trustworthiness of RS is a human-centered consideration. Therefore, apart from the technical side, evaluations on user perception and ethical side Wang et al. (2023) are equally vital. This therefore may require user studies to assess the subjective perceptions of fairness, robustness, explainability and even potential societal impacts.
@@
Causal discovery
Causal disentanglement
Conformity bias
An effective way to separate interaction data based on their primary causes
Relies on prior knowledge of causal models and is hard to scale to complex scenarios
[1]
Path discovery
Black-box mechanism
Straightforward to illustrate the process of recommendations
Potential risks of privacy leakage
[2]
Causal effect inference
Propensity reweighting
Position bias, selection bias
Proven to be statistically unbiased
Hard to accurately and robustly estimate the propensity weights
[3], [4], [5], [6]
Structure adjustment
Popularity bias, spurious correlations, bias amplification
Effective in deconfounding and estimating accurate causal effects
Relies on given causal models
[7], [8], [9], [10]
Uplift effect estimation
Uplift evaluation
Insightful to show the actual promotion brought by recommendations
Additional training data are required, which can be difficult to obtain in real-world applications
[11]
Counterfactual reasoning
Counterfactual augmentation
Population bias, poisoning attacks, spurious correlations
Excels in generating sufficient samples for training
Lack of assurance of the quality of generated samples
[12], [13]
Counterfactual inference
Attribute bias, popularity bias, noisy feedback, filter bubble bias, unexplainable results
A feasible way to explore causal effects of hypothetical interventions
Assumptions involved in computation may not always hold in practice
[14], [15], [16], [17], [18], [19], [20], [21]
[1]Zheng et al. (2021); [2]Xian et al. (2019); [3]Joachims et al. (2018); [4]Zhang et al. (2020b); [5]Li et al. (2023b); [6]Li et al. (2023a); [7]Zhang et al. (2021c); [8]Gupta et al. (2021); [9]He et al. (2023); [10]Wang et al. (2021b); [11]Sato et al. (2019); [12]Mu et al. (2022); [13]Xu et al. (2023a); [14]Li et al. (2021); [15]Wei et al. (2021); [16]Zhang et al. (2021b); [17]Wang et al. (2022b); [18]Gao et al. (2024); [19]Tan et al. (2021); [20]Chen et al. (2022); [21]Yu et al. (2023);
@@ 3 Causality-oriented TRS Methods
Recently, causal learning has stood out as an appealing solution to addressing trustworthiness challenges in the construction of TRS. We first present a taxonomy of existing CTRS methods, as illustrated in Figure 2, organized by different causal learning techniques upon which they are built. Then, we discuss their technical details and compare different classes of CTRS methods in Table 2, regarding their applicable challenges, pros and cons, and representative works.
@@ 3.1 Causal Discovery based Methods
Causal discovery is a crucial aspect of causal learning, focusing on identifying causal relationships from data. Unlike traditional data-driven methods that emphasize correlation, causal discovery based methods seek to uncover the causality among variables and formulate it in causal models. This is particularly valuable in RS, where understanding the causal relationships between each variable (e.g., users’ intention or items’ popularity) and the outcome (e.g., user-item interactions) is essential yet intractable. For different types of RS data, two representative causal discovery based methods have been researched: Causal Disentanglement based RS (CDRS) and Path Discovery based RS (PDRS).
@@ 3.1.1 Causal Disentanglement based RS (CDRS)
Interaction data are commonly influenced by various factors and underlying causes. For addressing fairness issues, it is essential to learn representations that remain unaffected by bias factors, such as user conformity. CDRS methods Zheng et al. (2021) address this by leveraging the colliding effect to distill cause-specific data from the original training dataset. This enables the learning of causally disentangled representations from the refined data, effectively reducing the impact of conformity bias, a common challenge in traditional RS. However, CDRS methods heavily rely on prior knowledge of causal relationships. Meanwhile, the complexity of disentanglement increases as more factors are involved.
@@ 3.1.2 Path Discovery based RS (PDRS)
Recent years have witnessed an upsurge in leveraging knowledge graphs Yang et al. (2022) for accurate recommendations. To explore the causality during knowledge graph reasoning and to enhance the understanding of the recommendation mechanism, PDRS Xian et al. (2019) trains a reinforcement learning agent to learn the navigation to suitable items for each user. An efficient path reasoning algorithm is devised to sample possible paths, which can be used as straightforward explanations of how recommendations are generated.
@@ 3.2 Causal effect inference based Methods
Based on the discovered or given causality in an RS, causal effect inference aims to understand and quantify the effect of one variable on another. Generally, causal effect inference based methods can be grouped into Propensity Reweighting based RS (PRRS), Structure Adjustment based RS (SARS) and Uplift Effect Estimation based RS (UEERS).
@@ 3.2.1 Propensity Reweighting based RS (PRRS)
Propensity reweighting is a set of fundamental techniques in causal effect inference, designed to adjust the biased observations in scenarios where a random experimental assignment is impractical. This makes it particularly useful in addressing position bias Joachims et al. (2018) and selection bias Li et al. (2023b) in RS. PRRS methods usually involve three key steps, i.e., propensity weight estimation, weight assignment and unbiased recommendation. In detail, propensity weight estimation aims to calculate the probability of an item receiving treatment (e.g., interaction). For example, in the problem of post-click conversion rate estimation Zhang et al. (2020b), the propensity weight is the click-through rate, which can be predicted through a logistic regression model. The inverse of the propensity weight can be then applied to the loss target, following the idea of Inverse Propensity Scoring (IPS) Hirano et al. (2003). Finally, these adjustments help mitigate biases, leading to more robust and fair recommendations. However, PRRS methods commonly face challenges with accuracy in weight estimation and high variance due to inverse weighting. Fortunately, recent advancements, such as IPS-V2 Li et al. (2023b) and doubly robust model Wang et al. (2019), have been developed to tackle these issues.
@@ 3.2.2 Structure Adjustment based RS (SARS)
In a given system where the underlying causal relationships are modeled by an SCM, structure adjustment emerges as a pivotal technique to accurately estimate causal effects, particularly by mitigating the influence of confounders. Recognizing that confounders are often the primary sources of various biases and noises, structure adjustment has gained great attention in building TRS. Depending on the variety of confounders, SARS methods usually adopt two types of adjustment tools. The first one is known as backdoor adjustment Zhang et al. (2021c), which leverages the backdoor criterion Pearl (2009) to identify and adjust observed confounders. Frontdoor adjustment Xu et al. (2023c), as another common technique, shows particular value in cases with unobserved confounders. Following the frontdoor criterion Pearl (2009), it finds a mediator to help in estimating the causal effects when direct adjustments for confounders are not possible.

### v2

@@ 2.3 Challenges in Representation Learning
Representation learning aims to learn informative user and item representations $\mathbf{h}_{u}$ and $\mathbf{h}_{i}$ from interaction data, bridging the stages of data preparation and recommendation generation. Although numerous advanced representation learning methods have been developed in RS, such as recurrent neural networks, graph neural networks Ma et al. (2020) and transformer-based models Sun et al. (2019), they commonly overlook potential risks regarding fairness and robustness.
@@ 2.3.1 Fairness Challenges in Representation Learning
A major fairness challenge is how to address conformity bias Zheng et al. (2021). This bias is closely related to popularity bias, but specifically focuses on the perspective of user intentions. In general, a user-item interaction is usually caused by two factors: users’ genuine interests $I_{\rm int}$ in the item and their conformity $I_{\rm con}$ to the item’s popularity, i.e., $I_{\rm int}\rightarrow r_{ui}\leftarrow I_{\rm con}$. An unbiased model should focus on learning users’ true interests rather than being misled by their conformity. However, it is intractable to clearly separate conformity intention from user representation, especially relying solely on statistical approaches. To tackle this challenge, Zheng et al. Zheng et al. (2021) introduced a causality-driven framework, which leverages the colliding effect Pearl (2009) to learn intention-disentangled item and user representations $\{\mathbf{h}_{u}^{\rm int},\mathbf{h}_{u}^{\rm con}\}$ and $\{\mathbf{h}_{i}^{\rm int},\mathbf{h}_{i}^{\rm con}\}$ from cause-specific data.
@@ 2.3.2 Robustness Challenges in Representation Learning
Underlying noise is a key robust challenge in representation learning. Here, we focus on two types: spurious correlations He et al. (2023) and noisy feedback Zhang et al. (2021b).
Spurious correlation refers to a correlation that is statistically strong but in fact is causally incorrect. To clarify this concept, consider a straightforward example. It is a well-established fact that an item’s quality influences both its price and the users’ preference for it. Commonly, one can observe that higher-quality items are more expensive and more appealing to users. However, if the factor of quality is neglected, an RS may falsely imply that user preference increases along with the price. Such spurious correlations can be largely misleading. Unfortunately, statistical methods, which typically rely on correlations or associations Luo et al. (2023) among samples, can hardly identify and eliminate such noise. In contrast, causality-oriented methods excel in tackling this issue by modeling and adjusting underlying causal relationships. For instance, recognizing that spurious correlations often stem from confounders, a study by Zhu et al. (2024) performs structure adjustment to mitigate corresponding confounder effects. Another representative research by Mu et al. (2022) introduces a counterfactual-based method to reduce spurious correlations by generating higher-quality counterfactual interactions.
Noisy feedback manifests in two forms: 1) positive interactions from users who ultimately turned out to be unsatisfied Wang et al. (2021a, c); and 2) delayed interactions from users who are actually satisfied Zhang et al. (2021b). Such noisy feedback limits the ability of an RS to capture users’ genuine preferences. To defend against this noise, Zhang et al. Zhang et al. (2021b) devised a counterfactual modifier in a reinforcement learning framework for streaming recommendation. It adopts importance sampling Yasui et al. (2020) and modifies the weight of reward in a counterfactual world to alleviate the impact of delayed feedback.
@@ 2.4 Challenges in Recommendation Generation
Upon obtaining representations from input data, RS models are devised to generate accurate suggestions $\hat{r}_{ui}=f(\mathbf{h}_{u},\mathbf{h}_{i}\mid\Theta)$ that are tailored to individual preferences. Despite the promising performance that modern RS models Lin et al. (2022) achieved, there still exist several challenges regarding fairness and explainability due to their intrinsic mechanisms.
@@ 2.4.1 Fairness Challenges in Recommendation Generation
Additionally, feedback loops in RS often reinforce users’ existing preferences, creating an echo chamber that isolates them from diverse contents, known as the filter bubble bias Gao et al. (2024b). To counter this challenge, a framework based on counterfactual inference Wang et al. (2022b) has been developed. This approach allows users to determine whether actions are needed to modify user embeddings through counterfactual modification. By doing so, it facilitates access to a wider and more diverse range of content, effectively helping users to break out of filter bubbles.
2.4.2 Explainability Challenges in Recommendation
The advancement of machine learning-based, especially deep learning-based techniques, has greatly enhanced the capability of RS models over the past decades. However, the inherent drawback of lacking interpretability has garnered significant concerns in recent years. With the introduction of complex black-box modules, it is challenging to explain how and why the recommendations are generated. This limitation not only impedes researchers in comprehensively understanding and improving models, but also undermines stakeholders’ trust in RS, especially in some safety-critical applications. Therefore, effective explainers are required to provide insightful explanations for both recommendation mechanisms and results.
Despite several attempts Zhang and Chen (2020) to generate explanations based on tags Vig et al. (2009) or visual attentions Chen et al. (2019), they are not convincing enough and often struggle to explain the mechanism of RS models. To bridge this gap, causality has been explored in RS to generate insightful explanations. For instance, a reinforcement learning model Xian et al. (2019) is designed to discover reasonable causal paths as explanations for how recommendations are generated. However, such path-based explanations also raise concerns regarding privacy leakage. Another representative work Tan et al. (2021) introduces a model-agnostic explainer based on counterfactual inference. It aims to find straightforward explanations by exploring how minimal changes $\Delta=\{\delta_{1},\delta_{2},\cdots,\delta_{R}\}$ in $R$ item features can alter recommendation outcomes $\hat{r}^{\Delta}_{ui}$. This counterfactual explanation not only interprets recommendation mechanisms but also has been proven to enhance recommendation performance by filtering out negative items Wang et al. (2024b).
@@ 2.5 Challenges in Evaluation
Moving beyond accuracy, TRS encompasses a variety of trustworthiness objectives, posing the necessity of specialized evaluation metrics. In particular, fairness metrics Li et al. (2023c) typically measure group or individual disparities. For instance, Wang et al. Wang et al. (2022b) employed an isolation index to measure the segregation across different user groups and to assess the severity of filter bubbles. In evaluating robustness, most works Li et al. (2023a, b) test RS accuracy under varying levels of robustness challenges. In evaluating explainability, user reviews Tan et al. (2021) are usually collected to assess the consistency with generated explanations. Moreover, based on the concepts of necessity and sufficiency, Tan et al. Tan et al. (2021) introduced two novel metrics to provide a thorough evaluation of explanation quality.
However, it should be pointed out that only a few attempts Sato et al. (2019) have been made to devise evaluations from a causal learning perspective, which is especially valuable when evaluation is required in unobservable cases, such as counterfactual scenarios. This remains an important yet underexplored direction and deserves greater research attention. Moreover, we keep in mind that the trustworthiness of RS is a human-centered concern. Therefore, beyond the technical aspects, evaluating user perception and ethical considerations Wang et al. (2023a) is equally vital. This may require user studies to assess the subjective perceptions of fairness, robustness, explainability, and even potential societal impacts.
@@
Causal discovery Causal disentanglement Conformity bias An effective way to separate interaction data based on their primary causes Relies on prior knowledge of causal models and is hard to scale to complex scenarios [1]
Path discovery Black-box mechanism Straightforward to illustrate the process of recommendations Potential risks of privacy leakage [2]
Causal effect inference Propensity reweighting Position bias, selection bias Proven to be statistically unbiased Hard to accurately and robustly estimate the propensity weights [3], [4], [5], [6]
Structure adjustment Popularity bias, spurious correlations, bias amplification Effective in deconfounding and estimating accurate causal effects Relies on given causal models [7], [8], [9], [10]
Uplift effect estimation Uplift evaluation Insightful to show the actual promotion brought by recommendations Additional training data are required, which can be difficult to obtain in real-world applications [11]
Counterfactual reasoning Counterfactual augmentation Population bias, poisoning attacks, spurious correlations Excels in generating sufficient samples for training Lack of assurance of the quality of generated samples [12], [13]
Counterfactual inference Attribute bias, popularity bias, noisy feedback, filter bubble bias, unexplainable results A feasible way to explore causal effects of hypothetical interventions Assumptions involved in computation may not always hold in practice [14], [15], [16], [17], [18], [19], [20], [21], [22]
[1]Zheng et al. (2021); [2]Xian et al. (2019); [3]Joachims et al. (2018); [4]Liu et al. (2023); [5]Li et al. (2023b); [6]Li et al. (2023a); [7]Zhang et al. (2021c); [8]Gupta et al. (2021); [9]He et al. (2023); [10]Wang et al. (2021b); [11]Sato et al. (2019); [12]Mu et al. (2022); [13]Xu and et al. (2023); [14]Li et al. (2021); [15]Wei et al. (2021); [16]Zhang et al. (2021b); [17]Wang et al. (2022b); [18]Gao et al. (2024b); [19]Tan et al. (2021); [20]Wang et al. (2024a); [21]Wang et al. (2024b); [22]Chen et al. (2024)
@@ 3 Causality-oriented TRS Methods
Upon reviewing the trustworthiness challenges at each stage of TRS and the necessity of causal solutions, we now examine various CTRS methods. First, we present a taxonomy of existing CTRS methods, as illustrated in Figure 2, categorized based on the causal learning techniques they employ. Then, we discuss their technical details and compare different classes of CTRS methods in Table 2, regarding their applicable challenges, pros and cons, and representative works.
@@ 3.1 Causal Discovery based Methods
Causal discovery, as a fundamental aspect of causal learning, identifies causal relationships from data. Unlike traditional methods that emphasize correlation Luo et al. (2023); Gao et al. (2024a), it uncovers the causality among variables and formulates causal models. This is particularly valuable in RS, where understanding the causal relationships between variables (e.g., users’ intention or items’ popularity) and outcomes (e.g., predicted interactions $\hat{r}_{ui}$) is essential yet intractable Wang et al. (2023b). Two key methods have been explored for RS data: Causal Disentanglement based RS (CDRS) and Path Discovery based RS (PDRS).
@@ 3.1.1 Causal Disentanglement based RS (CDRS)
Underlying factors (e.g., user intentions) drive various trustworthiness issues (e.g., popularity bias), yet they are difficult to disentangle, hindering the learning of representations unaffected by them. CDRS methods Zheng et al. (2021) tackle this by leveraging causality-based rules to distill cause-specific data from raw training data. For instance, based on the observation that both interest and conformity intentions affect interactions in a collider structure, $I_{\rm int}\rightarrow r_{ui}\leftarrow I_{\rm con}$, we can identify cause-specific data via the colliding effect Pearl (2009). This effect implies that knowing $r_{ui}$ makes $I_{\rm int}$ and $I_{\rm con}$ dependent; for example, interactions with lower $I_{\rm int}$ are more likely to be caused by higher $I_{\rm con}$. In this way, causally disentangled representations can be learned from the refined data, effectively reducing the impact of conformity bias. However, the flexibility and applicability of CDRS can be limited due to their heavy reliance on prior knowledge of causal models and the complexity of practical scenarios.
@@ 3.1.2 Path Discovery based RS (PDRS)
Recent years have witnessed an upsurge in leveraging knowledge graphs Yang et al. (2022) for accurate recommendations. To explore the causality during knowledge graph reasoning and to enhance the understanding of the recommendation mechanism, PDRS Xian et al. (2019) trains a reinforcement learning agent to learn the navigation to suitable items for each user. The explored path can be used as a straightforward explanation of how recommendations are generated.
@@ 3.2 Causal Effect Inference based Methods
Based on the discovered or prior causality in an RS, causal effect inference aims to understand and quantify the effect of one variable on another. Generally, causal effect inference based methods can be grouped into Propensity Reweighting based RS (PRRS), Structure Adjustment based RS (SARS), and Uplift Effect Estimation based RS (UEERS).
@@ 3.2.1 Propensity Reweighting based RS (PRRS)
Propensity reweighting is designed to adjust the biased observations when the random experimental assignment is impractical (e.g., the cases of position bias Joachims et al. (2018) and selection bias Li et al. (2023b)). PRRS methods usually involve three key steps, i.e., propensity weight estimation, weight assignment and unbiased recommendation. First, they estimate the propensity score $\hat{p}_{ui}$, which represents the probability of receiving treatment; for example, $\hat{p}_{ui}=P(o_{ui}=1)$ represents the likelihood of observing the true rating $r_{ui}$ in MNAR cases. This estimation is usually performed by statistical methods like Naive Bayes or neural networks such as multi-layer perceptron Liu et al. (2023). Then, $\hat{p}_{ui}$ is used to inversely weight the original prediction objective $\mathcal{L}$, e.g.,
[…]
Finally, the reweighted training process helps mitigate biases, leading to more robust and fair recommendations. However, PRRS methods commonly face challenges with accuracy in weight estimation and high variance due to inverse weighting. Fortunately, recent advancements, such as IPS-V2 Li et al. (2023b) and doubly robust model Zhang et al. (2020b), have been developed to tackle these issues.
@@ 3.2.2 Structure Adjustment based RS (SARS)
In a given system where the underlying causal relationships are modeled by an SCM, structure adjustment emerges as a pivotal technique to perform causal interventions, e.g., mitigating the influence of confounders. Recognizing that confounders are often the primary sources of biases and noise, various structure adjustment methods have gained great attention in building TRS, including backdoor adjustment Zhang et al. (2021c) and frontdoor adjustment Xu et al. (2023b). Backdoor adjustment utilizes the backdoor criterion Pearl (2009) to identify a set of observed variables $\mathbf{Z}$ that block all backdoor paths between the treatment $T$ and outcome $Y$. The causal effect is estimated by adjusting for $\mathbf{Z}$:
[…]
In RS, this method is applied to control for confounders like item popularity Zhang et al. (2021c), which affects both item and interaction. Frontdoor adjustment applies when confounders are unobservable, but there exists a mediator $\mathbf{M}$ that transmits the causal effect from $T$ to $Y$. Using the frontdoor criterion Pearl (2009), the causal effect is computed as:
[…]